A rectangular block of land is 240m x 180m in size. A well of uniform diameter 14m and 20m depth is dug in the middle of the block and the soil dug out is evenly spread over the remaining area of the land.
What will be the increase in the height of the block?

Answer by mananth(16946)

diameter of well = 14 m, radius = 7
depth = 20 m
Volume of mud dug out = pi*r^2*h
V= 3.14*7^2*20
V= 3077.2 m^3
Area of Land = 240* 180 = 43200 sq. m
Volume of the land after filling = L*W*H
dug out mud is the volume increase
3077.2 = 43200*H
H=3077.2/43200
H= 0.0712 m
0.0712 *100 = 7.12 cm
The height will increase by 7.12 cm
m.ananth@hotmail.ca